A stock’s return has the distribution below. Calculate the stock’s Sharpe ratio, expressed to three decimals. Assume the risk-free rate is 3%. (Note: it is easier to use your values from above and plug & chug.)

Demand for Products

Probability of Demand Occurring

Rate of Return if Demand Occurs

Weak

0.1

-30%

Below Average

0.1

-14%

Average

0.3

11%

Above Average

0.3

20%

Strong

0.2

45%
Answer #1

average expected return

=0.1*(-30%)+0.1*(-14%)+0.3*11%+0.3*20%+0.2*45%

=13.900%

standard deviation=(0.1*(-30%-13.90%)^2+0.1*(-14%-13.90%)^2+0.3*(11%-13.90%)^2+0.3*(20%-13.90%)^2+0.2*(45%-13.90%)^2)^(1/2)=21.86%

sharpe ratio=(13.90%-3%)/21.86%

=0.499